Two automobiles leave the same location traveling in opposite directions. The blue car travels x mph; the red car travels y mph. How long will it be before the cars are 200 miles apart, if the red car is traveling 10 mph slower?
Answer by stanbon(75887)

Two automobiles leave the same location traveling in opposite directions. The blue car travels x mph; the red car travels y mph. How long will it be before the cars are 200 miles apart, if the red car is traveling 10 mph slower?
----
Equations:
y = x-10 mph
---
The separation rate of the two cars is x+y or x+x-10 = 2x-10
----
time = distance/rate
---
time = 200/(2x-10) = 200/(2(x-5)) = 100/(x-5) hrs.
================================
Cheers,
Stan H.
============